what is the volume of a sphere with a radius of 2 cm?
What is the volume of a cone with 5cm height and 4cm width?
what is 1/2h(L1+L2)with L1:10.8, L2: 8.2, h:4?

Answer by still_happy2006(3)

I am only able to answer two of your questions:

1.) To find the volume of a sphere you use this formula:

[4/3] X [Pi] X [radius]^3 and is expressed in cubic units
[4/3] X [3.14] X [2 cm]^3 = [4/3] X [3.14] X [8]= 34.49 cubic centimeters.

2.) To find the volume of a cone you use this formula:
The volume of a cone is [1/3][PI][radius]^2[height]and is expressed in cubic units. Since you gave the full width the radius would be 2 cm.
[1/3][3.14][2cm]^2[5cm]= [1/3][3.14][4cm][5cm]= 20.93 cubic centimeters